# Weak Existence and Uniqueness for Super-Brownian Motion with Irregular Drift

### v1

@@ 6.1 An Artificial Branching Process
Let $Y$ be the process constructed in (6.5) with $Y_{0}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Then it follows that
$\sup_{n}\mathbb{E}\Big{[}\int_{0}^{T}\langle Y_{t},1\rangle dt\Big{]}\leq C(T),$ (6.17)
@@
$\begin{split}&\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{]}=\lim_{i\rightarrow\infty}\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s},1\rangle ds\mathbbm{1}_{t\leq\bar{T}_{i}}\Big{]}\\ &\leq\mathbb{E}[\int_{0}^{t}\langle\hat{Y}_{s},1\rangle ds].\end{split}.$
@@
$\begin{split}&\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{]}\leq\mathbb{E}\Big{[}\int_{0}^{t}\sum_{\alpha\in I_{s}}\langle Y_{s}^{(\alpha)},1\rangle ds\Big{]}\\ &\leq\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s}^{(1)},1\rangle ds\Big{]}+\mathbb{E}\Big{[}\sum_{\alpha\in I_{t}}\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds\Big{]}\\ &\leq\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds+C\mathbb{E}[I_{t}]\leq C(t),\end{split}$
@@
Let $Y_{0}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Then, for every $\gamma\in\mathbb{R}$, there exists $T_{0}>0$ such that for all $t\leq T_{0}$ we have
$\mathbb{E}\Big{[}\exp\Big{(}-\gamma\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{)}\Big{]}<C,$
Hereby, $T_{0}>0$ and $C>0$ only depend on $\langle\mu,1\rangle$ and $\nu((0,\infty])$.
@@
$\int_{0}^{t}\langle Y^{\alpha}_{s},1\rangle ds\leq\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds,\quad\forall\alpha\in I_{t},t\geq 0,$
@@
$\mathbb{E}\Big{[}\exp\Big{(}-\gamma\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{)}\Big{]}\leq\mathbb{E}\Big{[}\exp\Big{(}-\gamma|I_{t}|\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds-\gamma\int_{0}^{t}\langle V_{s}(\mu),1\rangle ds\Big{)}\Big{]}<C,$
for all $t\leq T_{0}$. This finishes the proof. ∎
@@ 6.2 Convergence of the Dual Process
Denote by $(Y_{t}^{n,m}(Y_{0}^{m}))_{t\geq 0}$ the process as constructed in (6.5) with initial condition $Y_{0}^{m}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$ and $n\in\mathbb{N}\cup\{\infty\}$, $m\in\mathbb{R}_{+}\cup\{\infty\}$. We assume that $Y_{0}^{n,m}=Y_{0}^{m}$, $\langle Y_{0}^{m},1\rangle>0$ and that, for $n\in\mathbb{N}$, $\nu$ is replaced by
$\nu_{n}=\nu^{1}_{|[0,n]}+\nu^{2}_{|[0,n]}+d_{1}\delta_{n}+d_{2}\delta_{n}=\nu^{1}_{n}+\nu^{2}_{n}+d_{1}\delta_{n}+d_{2}\delta_{n}.$
Here, $d_{1}$, $d_{2}$ and $a$ are defined as in (A). For $i\geq 1$, we denote by $\bar{T}_{n,i}^{m}$ the jump times, by $T_{n,i}=\bar{T}_{n,i}^{m}-\bar{T}_{n,i-1}^{m}$ the time in between jumps, by $Z_{n,i}^{m}$ the jump sizes and by $U_{n,i}^{m}$ the jump positions defined as in (6.2)-(6.4). Moreover, we denote by $M_{n,i}^{m}$ the random variables defined as in (6.1). Finally, we shall denote by $(S_{n,k}^{m})_{k\geq 1}$ the sequence independent of $\exp(1)$-distributed random variables from the construction of the process $Y^{n,m}$ in the beginning of Section 6.
@@
for $t\in[0,T_{0}]$, where $T_{0}$ is chosen so that (6.8) holds with $\gamma=2a$ and where $X_{0}$ is a bounded and continuous function on $\mathbb{R}$.
Note, that $T_{0}$ can be chosen independent of $n,m$ since
@@
$|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{x}},s)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|\leq|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s+\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|+|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s-\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|$
@@
$\begin{split}&\eta(\delta)\leq\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,1\rangle dt\\ &\leq 2\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt\\ &\qquad+\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{\bar{t}}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,\mathbbm{1}_{D_{R}}\rangle dt\\ &\qquad+2k\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,\infty),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt.\end{split}$ (6.25)
@@
$\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt+k\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,0),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt\quad<\tilde{\epsilon}$.
@@
    $\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,m}}\exp\Big{(}-\langle Y^{\infty,m}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y^{\infty,m}_{s},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y^{\infty,\infty}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\Big{)}\Big{]}\end{split}$
@@
We only show (i), since (ii) and (iii) follow along similar lines. Let $T_{0}>0$ be such that (6.8) is satisfied with $\gamma=2a$. Fix arbitrary $t>0$ and let $\epsilon>0$. By (6.11), there exists a sequence of random variables $(\bar{R}_{i})_{i\geq 1}$ such that
@@
Let $T_{0}>0$ be such that (6.8) holds for $\gamma=2a$ and let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with initial condition $X_{0}$ satisfying (1.5). Furthermore, let $(g_{n})_{n\geq 1}$ be a sequence of measurable bounded functions on $\mathbb{R}$, that converge boundedly pointwise to a bounded, measurable function $g$ on $\mathbb{R}$. Then it follows for all $T\leq T_{0}$ that
@@
Fix $m\in\mathbb{N}$. To lighten notation, we suppress the subscript $m$ for the rest of the proof. Note that by Lemma 6.2 and (6.13) we have
$\sup_{n}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{4}\Big{]}\leq C\sup_{n}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n},1\rangle ds\Big{)}^{4}\Big{]}\leq C\mathbb{E}[I_{T}^{4}]<\infty$
Furthermore, by Lemma 6.8 it follows that
$\begin{split}&(U_{n,1},U_{n,2},\ldots,T_{n,1},T_{n_{2}},\ldots,Z_{n,1},Z_{n,2},\ldots)\\ &\qquad\Rightarrow(U_{\infty,1},U_{\infty,2},\ldots,T_{\infty,1},T_{\infty,2},\ldots,Z_{\infty,1},Z_{\infty,2},\ldots)\end{split}$ (6.47)
and thus by the Skorohod representation theorem there exists a probability space such that the convergence (6.47) holds almost surely. We now show that almost surely
$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds=0.$
First, we can use a similar argument to the proof of Lemma 6.8 to see that for almost all $\omega$ there exists $M(\omega)\in\mathbb{N}$ such that
$\sup_{n}\bar{T}(\omega)_{n,M(\omega)+1}>T\quad,M(\omega)<\infty.$
Furthermore, note that there exists $D(\omega)>0$ such that almost surely
@@
for all $k\leq M$, since $(U_{n,k})_{n\geq 1}$ converges almost surely for each fixed $k$. Let $\epsilon>0$ be arbitrary. Then, by using Lemma 3.2 and Proposition 3.6 we get the following. There exists $\bar{D}(\omega)>D(\omega)$ such that
$\int_{0}^{T}\int_{B(0,\bar{D})^{c}}Y^{n}_{s-}(x)dxds\leq M\int_{0}^{T}\int_{B(0,\bar{D})^{c}}\big{(}\mathcal{W}(D,0)\big{)}(x)dxds\leq\epsilon,$
with $\bar{D}(\omega)<\infty$ almost surely. Furthermore, there exists $\delta_{1}(\omega)>0$ such that
$M\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds<\epsilon.$
@@
almost surely. Thus, we get
$\begin{split}&\int_{0}^{T}\langle Y_{s-}^{n},|h_{n}(X_{T-s})-h(X_{T-s})|\rangle ds\\ &\leq C\epsilon+\sum_{k=1}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq C\epsilon+\int_{0}^{T}\int_{B(0,\bar{D})}\big{(}V_{s}(\phi+m\delta_{x_{0}})\big{)}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\quad+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds,\end{split}$
@@
$\begin{split}&\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq CM\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n+\delta_{1}}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|h_{n}(X_{T-s}(x))-h(X_{T-s}(x))|dxds\\ &\leq C\epsilon+M\big{(}\sup_{s\in(\delta_{1},T),x\in B(0,\bar{D}+D)}\mathcal{W}_{s}(0,0)(x)\big{)}\int_{0}^{T}\int_{B(0,\bar{D})}|h_{n}(X_{T-s}(x))-h(X_{T-s}(x))|dxds\\ &\leq C\epsilon+K\int_{0}^{T}\int_{B(0,\bar{D})}|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\end{split}$
@@
$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\leq 2C\epsilon$
almost surely. Since $\epsilon>0$ was arbitrary, this proves that
@@
almost surely. Thus, by uniform integrability, we get
$\begin{split}&\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{2}]\\ &=\mathbb{E}\Big{[}\Big{(}\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{2}\Big{]}=0.\end{split}$

### v2

@@ 6.1 An Artificial Branching Process
Let $Y$ be the process constructed in (6.6) with $Y_{0}^{m}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Then it follows for all $T>0$ that
$\sup_{m}\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{T}\langle Y_{t},1\rangle dt\Big]\leq C(T),$ (6.21)
@@
$\begin{split}&\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\langle Y_{s},1\rangle ds\Big]=\lim_{i\rightarrow\infty}\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\langle Y_{s},1\rangle ds\mathbbm{1}_{t\leq\bar{T}_{i}}\Big]\\ &\leq\mathbb{E}_{Y_{0}^{m}}[\int_{0}^{t}\langle\hat{Y}_{s},1\rangle ds].\end{split}.$
@@
$\begin{split}&\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\langle Y_{s},1\rangle ds\Big]\leq\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\sum_{\alpha\in I_{s}}\langle Y_{s}^{(\alpha)},1\rangle ds\Big]\\ &\leq\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\langle Y_{s}^{(1)},1\rangle ds\Big]+\mathbb{E}\Big[\sum_{\alpha\in I_{t}}\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds\Big]\\ &\leq\int_{0}^{t}\langle V_{s}(Y_{0}^{m}),1\rangle ds+C(t)\mathbb{E}[I_{t}]\leq C(t),\end{split}$
@@
Let $Y_{0}^{m}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Let $Y(\nu)$ be the process constructed above with measure $\nu$. Then, for every $\gamma\in\mathbb{R}$ and $K>0$, there exists $T_{0}>0$ such that for all $t\leq T_{0}$ we have
$\sup_{\nu:\nu((0,\infty])\leq K}\sup_{m}\mathbb{E}_{Y_{0}^{m}}\Big[\exp\Big(-\gamma\int_{0}^{t}\langle Y_{s}(\nu),1\rangle ds\Big)\Big]<C,$ (6.22)
Hereby, $T_{0}>0$ and $C>0$ only depend on $\langle\mu,1\rangle$ and $K$.
@@
$\int_{0}^{t}\langle Y^{\alpha}_{s}(\nu),1\rangle ds\leq\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds,\quad\forall\alpha\in I_{t},t\geq 0,$
@@
$\mathbb{E}_{Y_{0}^{m}}\Big[\exp\Big(-\gamma\int_{0}^{t}\langle Y_{s}(\nu),1\rangle ds\Big)\Big]\leq\mathbb{E}\Big[\exp\Big(-\gamma|I_{t}^{\nu}|\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds-\gamma\int_{0}^{t}\langle V_{s}(\mu),1\rangle ds\Big)\Big]<C,$
for all $t\leq T_{0}$, $m\geq 1$ and $\nu$ with $\nu((0,\infty])\leq K$. This finishes the proof. ∎
@@ 6.2 Convergence of the Dual Process
Let $\nu$ be as in (6.1). For $n\in\mathbbm{N}\cup\{\infty\}$ and $m\in\mathbb{R}_{+}\cup\{\infty\}$ denote by $(Y_{t}^{n,m}(Y_{0}^{m}))_{t\geq 0}$ the process as constructed in (6.6) with initial condition $Y_{0}^{m}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$ and the measure $\nu$ is replaced by
$\nu_{n}=\nu^{1}_{|[0,n]}+\nu^{2}_{|[0,n]}+d_{1}\delta_{n}+d_{2}\delta_{n}=\nu^{1}_{n}+\nu^{2}_{n}+d_{1}\delta_{n}+d_{2}\delta_{n},$
for $n\in\mathbb{N}$. Here, $d_{1}$, $d_{2}$ and $a$ are defined as in (A) and for $n=\infty$ we set $\nu_{\infty}\equiv\nu$. We also assume that $\langle Y_{0}^{m},1\rangle>0$. For $i\geq 1$, we denote by $\bar{T}_{n,i}^{m}$ the jump times, by $T_{n,i}=\bar{T}_{n,i}^{m}-\bar{T}_{n,i-1}^{m}$ the time in between jumps, by $Z_{n,i}^{m}$ the jump sizes and by $U_{n,i}^{m}$ the jump positions defined as in (6.3)-(6.5). Moreover, we denote by $M_{n,i}^{m}$ the random variables defined as in (6.2). Finally, we shall denote by $(S_{n,k}^{m})_{k\geq 1}$ the sequence independent of $\exp(1)$-distributed random variables from the construction of the process $Y^{n,m}$ in the beginning of Section 6. Let $X_{0}$ be a bounded, continuous and non-neagtive function on $\mathbb{R}$.
@@
for $t\in[0,T_{0}]$. Here $T_{0}$ is chosen as in Lemma 6.5 with $\gamma=2a$ and $K=\nu((0,\infty])$.
Recall that by Lemma 6.5, $T_{0}$ can be chosen independent of $n,m$ since
@@
$|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{x}},s)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|\leq|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s+\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|+|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s-\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|.$
@@
$\begin{split}&\eta(\delta)\leq\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,1\rangle dt\\ &\leq 2\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt\\ &\qquad+\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{\bar{t}}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,\mathbbm{1}_{D_{R}}\rangle dt\\ &\qquad+2(k+1)\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,\infty),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt.\end{split}$ (6.30)
@@
$\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt+(k+1)\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,0),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt\quad<\tilde{\epsilon}$.
@@
    $\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big[(-1)^{J_{t}^{\infty,m}}\exp\Big(-\langle Y^{\infty,m}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y^{\infty,m}_{s},1\rangle ds\Big)\Big]\\ &=\mathbb{E}\Big[(-1)^{J_{t}^{\infty,\infty}}\exp\Big(-\langle Y^{\infty,\infty}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\Big)\Big],\end{split}$
@@
We only show (i), since (ii) and (iii) follow along similar lines. Let $T_{0}>0$ be such that (6.10) is satisfied with $\gamma=-2a$. Fix arbitrary $t>0$ and let $\epsilon>0$. By (6.14), there exists a sequence of random variables $(\bar{R}_{i})_{i\geq 1}$ such that
@@
Let $T_{0}>0$ be such that (6.22) in Lemma 6.5 holds for $\gamma=2a$ and $K=\nu((0,\infty])$. Furthermore, let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with initial condition $X_{0}$ satisfying (1.5). Furthermore, let $(g_{n})_{n\geq 1}$ be a sequence of measurable bounded functions on $\mathbb{R}$, that converge boundedly pointwise to a bounded, measurable function $g$ on $\mathbb{R}$. Then it follows for all $T\leq T_{0}$ that
@@
Fix $m\in\mathbb{N}$. To lighten notation, we suppress the superscript $m$ for the rest of the proof. Moreover, fix arbitrary $T\leq T_{0}$. Note that by Lemma 6.5 with $K=\nu((0,\infty])$ we have
$\sup_{n}\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big)^{4}\Big]\leq c\sup_{n}\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s-}^{n},1\rangle ds\Big)^{4}\Big]<\infty.$ (6.52)
Furthermore, we can use a similar argument to the proof of Lemma 6.9 to see that there exists $M\in\mathbb{N}$ such that
$\mathbb{P}(\inf_{n}\bar{T}_{n,M}\leq T)<\epsilon.$
and by Lemma 6.8 it follows that
$\begin{split}&(U_{n,1},\ldots,U_{n,M},T_{n,1},\ldots,T_{n,M},Z_{n,1},\ldots,Z_{n,M})\\ &\qquad\Rightarrow(U_{\infty,1},\ldots,U_{\infty,M},T_{\infty,1},\ldots,T_{\infty,M},Z_{\infty,1},\ldots,Z_{\infty,M})\end{split}$ (6.53)
and thus by the Skorohod representation theorem there exists a probability space such that the convergence (6.53) holds almost surely. We now show that
$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds=0$
for almost all $\omega$ in the event
$A_{\epsilon}=\{\inf_{n}\bar{T}_{n,M}>T\}.$
To this end, note that there exists $D(\omega)>0$ for $\omega\in A_{\epsilon}$ such that almost surely on $A_{\epsilon}$,
@@
for all $k\leq M$, since $(U_{n,k})_{n\geq 1}$ converges almost surely for each fixed $k$. Let $\tilde{\epsilon}>0$ be arbitrary. Then, by using Lemma 3.2 and Proposition 3.6 we get the following. There exists $\bar{D}(\omega)>D(\omega)$ such that for $\omega\in A_{\epsilon}$ we have
$\int_{0}^{T}\int_{B(0,\bar{D})^{c}}Y^{n}_{s-}(x)dxds\leq M\int_{0}^{T}\int_{B(0,\bar{D})^{c}}\big(\mathcal{W}(D,0)\big)(x)dxds\leq\tilde{\epsilon}$,
with $\bar{D}(\omega)<\infty$ a.e. on $A_{\epsilon}$. We continue to work with $\omega\in A_{\epsilon}$. There exists $\delta_{1}>0$ such that
$M\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds<\tilde{\epsilon}$.
@@
almost surely on $A_{\epsilon}$. Thus, we get
$\begin{split}&\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\\ &\leq C\tilde{\epsilon}+\sum_{k=1}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq C\tilde{\epsilon}+\int_{0}^{T}\int_{B(0,\bar{D})}\big(V_{s}(\phi+m\delta_{x_{0}})\big)(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\quad+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds,\end{split}$
@@
$\begin{split}&\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq CM\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n+\delta_{1}}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq C\tilde{\epsilon}+M\big(\sup_{s\in(\delta_{1},T),x\in B(0,\bar{D}+D)}\mathcal{W}_{s}(0,0)(x)\big)\int_{0}^{T}\int_{B(0,\bar{D})}|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq C\tilde{\epsilon}+K\int_{0}^{T}\int_{B(0,\bar{D})}|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\end{split}$
@@
$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\leq 2C\tilde{\epsilon}$
almost surely on $A_{\epsilon}$. Since $\tilde{\epsilon}>0$ was arbitrary, this proves that
@@
for almost all $\omega\in A_{\epsilon}$. Thus, by uniform integrability (see (6.52)), we get
$\begin{split}\lim_{n\rightarrow\infty}\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big)^{2}\mathbbm{1}_{A_{\epsilon}}]=0.\end{split}$
Finally, by Hölder’s inequality and (6.52), we get
$\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big)^{2}\mathbbm{1}_{(A_{\epsilon})^{c}}]\leq c\epsilon\sqrt{\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s}^{n},1\rangle ds\Big)^{4}\Big]}\leq C\epsilon,$
and since $\epsilon>0$ was arbitrary, this finishes the proof. ∎